# Both Stellar Mass and Gravitational Potential Shape the Gas-Phase Metallicity

### v1

@@ Abstract
The relation between metallicity and galaxy mass (the so-called mass-metallicity relation) is the strongest and most prominent among scaling relations between chemical enrichment and galactic properties. However, it is unclear whether this relation primarily traces metal retention or the integrated production of metals, as past studies have obtained contrasting results. We investigate this issue through an extensive Random Forest and Partial Correlations analysis of spectral cubes of 4,500 galaxies from the MaNGA survey. We find that stellar mass ($\rm M_{*}$) and baryonic gravitational potential ($\rm\Phi_{*}=M_{*}/R_{e}$) are the two most important quantities determining gas metallicity in galaxies. However, their relative roles strongly depend on the galactocentric radius – the metallicity within 1 $\rm R_{e}$ depends primarily on the stellar mass, while the metallicity at radii beyond 1.5 $\rm R_{e}$ depends primarily on the gravitational potential. This finding can be interpreted in terms of metals in the central region ($\rm R\leq 1~R_{e}$) being mostly bound, regardless of the global gravitational potential and, therefore, the metallicity is determined primarily by the cumulative production of metals (hence the integrated star formation history, i.e. $\rm M_{*}$); by contrast, in the galactic peripheries the retention of metals depends more critically on the gravitational potential, hence the stronger dependence of the metallicity on $\rm\Phi_{*}$ at large radii. Our finding reconciles apparent discrepancies between previous results. Finally, we find that the Star Formation Rate is the third most important parameter (after $\rm M_{*}$ and $\rm\Phi_{*}$) in determining the metallicity, as expected from the Fundamental Metallicity Relation.
@@
pubyear: 2025pagerange: Both Stellar Mass and Gravitational Potential Shape the Gas-Phase MetallicityG
@@ 1 Introduction
Galaxy evolution is governed by the regulation of gas content, metallicity, and stellar mass, shaped by gas inflows from the IGM, feedback-driven outflows, and internal recycling (Madau & Dickinson, 2014; Lilly et al., 2013; Péroux & Howk, 2020). Metallicity is linked to the time-integrated production of chemical elements; hence, it reflects a galaxy’s star formation history (Maiolino & Mannucci, 2019). Metallicity is also more broadly connected to the baryon cycle within and around galaxies. Indeed, inflows dilute ISM metals while fueling star formation, whereas stellar winds and supernovae enrich the ISM and CGM (Tumlinson et al., 2017; Lilly et al., 2013; Peng & Maiolino, 2014b). AGN feedback further influences the ISM and CGM by heating or expelling gas, suppressing star formation (Fabian, 2012; Bourne & Yang, 2023). On the other hand, AGN feedback can also trigger star formation within galactic outflows (Maiolino et al., 2017). Dust also affects metal depletion, and over time, gas cycles through phases of enrichment and redistribution. Given all these interconnected aspects between chemical enrichment and galaxy evolution, a strong relationship is expected between metallicity and galactic properties.
@@
The first, and most widely studied relation is between gas-phase metallicity and stellar mass, the so-called mass-metallicity relation (MZR; see e.g., Tremonti et al., 2004; Kewley & Ellison, 2008; Sánchez et al., 2017; Maiolino & Mannucci, 2019, for a review). This correlation was first defined as a metallicity versus luminosity relation by Lequeux et al. (1979), but later shown to be more directly connected with the galaxy mass. Specifically, an increase in stellar mass correlates with an increase in metallicity until the relation flattens out at high stellar masses ($\mathrm{\gtrsim 10^{10}M_{\odot}}$). This relation has been studied for the local Universe and redshifts up to $z\sim 10$ (Maiolino et al., 2008; Lamareille et al., 2009; Mannucci et al., 2010; Maier et al., 2014; Cresci et al., 2019; Curti et al., 2020, 2023; Langeroodi et al., 2023; Nakajima et al., 2023). The MZR has also been studied at spatially resolved scales (Rosales-Ortega et al., 2012; Sánchez et al., 2013; Sánchez Almeida & Sánchez-Menguiano, 2019; Baker et al., 2022b). This relation has been observed to evolve with redshift: galaxies in the early universe are metal-poorer than local galaxies (Tremonti et al., 2004; Savaglio et al., 2005; Maiolino et al., 2008; Gao et al., 2018; Sanders et al., 2021). Generally, this is attributed to galaxies in the local universe having had more time to assemble and produce metals than at high redshifts.
@@
The second most widely studied relation is the metallicity’s anti-correlation with the star formation rate (SFR), which, together with the stellar mass, gives rise to the so-called Fundamental Metallicity Relation (FMR; Mannucci et al., 2010; Lara-López et al., 2010; Cresci et al., 2019; Curti et al., 2020; Pistis et al., 2024; Curti et al., 2024; Sarkar et al., 2025). This relation has also been observed to exist at spatially resolved scales (Baker et al., 2022b; Koller et al., 2024). Other significant relations of the gas-phase metallicity have been found with the stellar age (Lian et al., 2015; Sánchez-Menguiano et al., 2020), rotational velocity (Garnett, 2002; Pilyugin et al., 2004; Hughes et al., 2013), galaxy size (Ellison et al., 2008; Sánchez Almeida & Dalla Vecchia, 2018), velocity dispersion (Li et al., 2018; Cappellari, 2023), and gas mass fraction (Bothwell et al., 2013; Barrera-Ballesteros et al., 2018; Brown et al., 2018).
@@
Nevertheless, stellar mass remains the strongest contributor to defining metallicity among those, and the origin of the MZR has long been debated (Maiolino & Mannucci, 2019). A common explanation is that massive galaxies inhabit a deeper gravitational potential that is more efficient at retaining metals against outflows compared to lower mass galaxies (Tremonti et al., 2004; Tumlinson et al., 2011; Chisholm et al., 2018). In this sense, both the dynamical mass ($\mathrm{M_{dyn}}$) and gravitational potential, defined as $\mathrm{\Phi_{dyn}=M_{dyn}/R_{e}}$, would form a stronger correlation with the metallicity than the stellar mass would. Dynamical measurements are much rarer than photometric or spectroscopic estimates; thus, a common approximation of the gravitational potential is $\mathrm{\Phi_{*}=M_{*}/R_{e}}$ (D’Eugenio et al., 2018; Cappellari, 2023).
@@
D’Eugenio et al. (2018) investigated the correlation between the metallicity and the stellar mass $\mathrm{M_{*}}$, gravitational potential $\mathrm{\Phi_{*}}$, and surface mass density $\mathrm{\Sigma_{*}=M_{*}/R_{e}}$ for a volume-limited sample from SDSS DR7. They found that the metallicity more tightly correlates with $\mathrm{\Phi_{*}}$ rather than $\mathrm{M_{*}}$ or $\mathrm{\Sigma_{*}}$.
@@
Sánchez-Menguiano et al. (2024a) used data from the MaNGA pyPipe3D value-added catalogue (VAC) of more than 3000 galaxies to simultaneously investigate the effect of 148 parameters on the gas-phase metallicity via random forest regression. Their results confirm that $\mathrm{\Phi_{*}}$, as a proxy for the baryonic gravitation potential, is the dominating property in determining the gas-phase metallicity. They also find that the photometric mass is more important in defining the metallicity than the stellar mass measured via spectroscopy.
@@
A tight correlation with $\mathrm{\Phi_{*}}$ has also been found for the stellar metallicity (Barone et al., 2018; Vaughan et al., 2022; Cappellari, 2023; Sánchez-Menguiano et al., 2024b).
@@
On the contrary, an investigation done by Baker & Maiolino (2023) found that the stellar mass $\mathrm{M_{*}}$, not $\mathrm{\Phi_{*}}$ or the dynamical measurements $\mathrm{M_{dyn}}$ and $\mathrm{\Phi_{dyn}}$, is the dominating factor in determining the gas-phase metallicity. Their study focused on Partial Correlation Coefficient (PCC) analysis and random forest regression (Bluck et al., 2020a; Baker et al., 2022a) of roughly 1000 galaxies from the MaNGA survey.
@@
Throughout this paper, we assume a Kroupa (2001) initial mass function (IMF) and adopt a flat $\mathrm{\Lambda CDM}$ cosmology with $\mathrm{H_{0}=70.0kms^{-1}Mpc^{-1}}$, $\mathrm{\Omega_{m}=0.3}$, and $\mathrm{\Omega_{\lambda}=0.7}$.
@@ 2.1 MaNGA
We utilise publicly available integral field spectroscopic (IFS) data from MaNGA (Mapping Nearby Galaxies at Apache Point Observatory, Bundy et al. 2015), part of the fourth-generation Sloan Digital Sky Survey (SDSS-IV, Blanton et al., 2017). We use the final and fully complete data release (DR) 17, released in 2022, for 10,000 galaxies (Abdurro’uf et al., 2022; Sánchez et al., 2022).
@@
We also use dynamical masses $\mathrm{M_{dyn}}$ and effective velocity dispersions, $\mathrm{\sigma_{e}}$, within elliptical half-light isophotes provided by Zhu et al. (2023), which were computed via the axisymmetric Jeans Anisotropic Modelling (JAM, Cappellari, 2008, 2020).
@@ 2.2 SDSS
Our second dataset consists of data from the Sloan Digital Sky Survey (SDSS) DR7 (Abazajian et al., 2009), which includes spectroscopy for around 930,000 galaxies. We get stellar masses and SFRs from the MPA-JHU catalogue [^2] https://wwwmpa.mpa-garching.mpg.de/SDSS/DR7/ (Kauffmann et al., 2003; Brinchmann et al., 2004; Salim et al., 2007) and use the provided emission line fluxes for computing gas-phase metallicities. We match this data to a catalogue containing morphological parameters (Simard et al., 2011).
@@ 2.3 Sample selection
Regarding the SDSS data, we employ a signal-to-noise ratio (SNR) cut $\rm S/N>3$ for our lines of interest to ensure accurate emission line detection for calculating gas-phase metallicities. We also limit the sample by excluding highly inclined galaxies by applying a minor-to-major axis ratio cut of $\mathrm{b/a\geq 0.35}$. Furthermore, due to the small aperture of the SDSS observations, some galaxies which are at a very low redshift might only be probed within their most central region. To reduce aperture effects, we restrict our sample to galaxies for which the spectroscopic fibre radius covers more than $0.5\mathrm{R_{e}}$. The calibrations used for computing gas-phase metallicities are based on star-forming regions; thus, we select star-forming galaxies based on the $\mathrm{[NII]/H\alpha}$ vs. $\mathrm{[OIII]/H\beta}$ Baldwin-Philips-Terlevich (BPT) diagnostic diagram (Baldwin et al., 1981), using the empirical line by Kauffmann et al. (2003), which separates star-forming galaxies from AGN/LINER galaxies. Our final SDSS sample consists of 121,880 star-forming galaxies.
@@ 2.4 Derivation of Physical Quantities
Stellar masses per pixel, integrated total stellar masses and masses within $\mathrm{1R_{e}}$ are obtained from pyPipe3D (Sánchez et al., 2022). We obtain galactocentric distances for each spaxel in units of the effective radius from DAP (Belfiore et al., 2019a; Westfall et al., 2019). Via the galactocentric distances, we compute total masses and masses within $\mathrm{1R_{e}}$ by taking the sum of all spaxels.
Emission lines are corrected for extinction using the $\mathrm{H\alpha/H\beta}$ Balmer decrement, assuming an intrinsic value of $2.86$ (Osterbrock & Ferland, 2006), via the Calzetti et al. (2000) dust attenuation law with $\mathrm{R_{v}=3.1}$.
Star formation rates for each spaxel in the MaNGA data are computed via the conversion law from Kennicutt & Evans (2012) using the $\mathrm{H\alpha}$ emission line:
@@
We obtain a total SFR within $1\mathrm{R_{e}}$ by taking the sum of SFRs of the spaxels which fulfil our $\rm S/N$ and BPT diagram requirements. The global catalogue of pyPipe3D data available online [^3] https://data.sdss.org/datamodel/files/MANGA_PIPE3D/MANGADRP_VER/PIPE3D_VER/SDSS17Pipe3D.html also provides integrated SFRs using only the spaxels compatible with star formation. The global values are computed using the Kennicutt (1998) calibration for a Salpeter IMF. To convert to the new calibration using a Kroupa IMF, we use the conversion given in Kennicutt & Evans (2012):
@@
Table 1: Diagnostics used for computing gas-phase metallicities as presented in Curti et al. (2020).
@@
To obtain a singular value for the gas-phase metallicities, we compute the metallicities per spaxel using nine different strong-line calibrations (see Table 1) and take the average over all spaxels from our selection within each galaxy. The diagnostics $\mathrm{R23,R2,R3,O32,N2,S2,O3S2,O3N2,}$ and $\mathrm{R3S2}$ are used in combination with the empirical calibrations presented in Curti et al. (2017, 2020). Using several diagnostics simultaneously decreases the risk of having degeneracies associated with diagnostics that are double values. It also allows for integrating information from several emission lines that may differ across metallicity regimes, and it allows us to constrain the resulting abundances more precisely (see Curti et al. 2020 for more information). We take the average metallicity value across all selected spaxels within $1\mathrm{R_{e}}$ to obtain a single value for the gas-phase metallicity per galaxy. We also repeated our analysis using the median and found no significant difference in the results.

### v2

@@ Abstract
The relation between metallicity and galaxy mass (the so-called mass-metallicity relation) is the strongest and most prominent among scaling relations between chemical enrichment and galactic properties. However, it is unclear whether this relation primarily traces metal retention or the integrated production of metals, as past studies have obtained contrasting results. We investigate this issue through an extensive Random Forest and Partial Correlations analysis of spectral cubes of 4,500 galaxies from the MaNGA survey. We find that stellar mass ($\rm M_{*}$) and baryonic gravitational potential ($\rm\Phi_{*}=M_{*}/R_{e}$) are the two most important quantities determining gas metallicity in galaxies. However, their relative roles strongly depend on the galactocentric radius – the metallicity within 0.7 $\rm R_{e}$ depends primarily on the stellar mass, while the metallicity at radii beyond 0.9 $\rm R_{e}$ depends primarily on the gravitational potential. This finding can be interpreted in terms of metals in the central region ($\rm R\leq 0.7~R_{e}$) being mostly bound, regardless of the global gravitational potential and, therefore, the metallicity is determined primarily by the cumulative production of metals (hence the integrated star formation history, i.e. $\rm M_{*}$); by contrast, in the galactic peripheries the retention of metals depends more critically on the gravitational potential, hence the stronger dependence of the metallicity on $\rm\Phi_{*}$ at large radii. Our finding reconciles apparent discrepancies between previous results. Finally, we find that the Star Formation Rate is the third most important parameter (after $\rm M_{*}$ and $\rm\Phi_{*}$) in determining the metallicity, as expected from the Fundamental Metallicity Relation.
@@
pubyear: 2025pagerange: Both Stellar Mass and Gravitational Potential Shape the Gas-Phase MetallicityF
@@ 1 Introduction
Galaxy evolution is governed by the regulation of gas content, metallicity, and stellar mass, shaped by gas inflows from the IGM, feedback-driven outflows, and internal recycling (madau_cosmic_2014; lilly_gas_2013; peroux_cosmic_2020). Metallicity is linked to the time-integrated production of chemical elements; hence, it reflects a galaxy’s star formation history (maiolino_re_2019). Metallicity is also more broadly connected to the baryon cycle within and around galaxies. Indeed, inflows dilute ISM metals while fueling star formation, whereas stellar winds and supernovae enrich the ISM and CGM (tumlinson_circumgalactic_2017; lilly_gas_2013; peng_haloes_2014). AGN feedback further influences the ISM and CGM by heating or expelling gas, suppressing star formation (fabian_observational_2012; bourne_recent_2023). On the other hand, AGN feedback can also trigger star formation within galactic outflows (maiolino_star_2017). Dust also affects metal depletion, and over time, gas cycles through phases of enrichment and redistribution. Given all these interconnected aspects between chemical enrichment and galaxy evolution, a strong relationship is expected between metallicity and galactic properties.
@@
The first, and most widely studied relation is between gas-phase metallicity and stellar mass, the so-called mass-metallicity relation (MZR; see e.g., tremonti_origin_2004; kewley_metallicity_2008; sanchez_mass-metallicity_2017; maiolino_re_2019, for a review). This correlation was first defined as a metallicity versus luminosity relation by lequeux_chemical_1979, but later shown to be more directly connected with the galaxy mass. Specifically, an increase in stellar mass correlates with an increase in metallicity until the relation flattens out at high stellar masses ($\mathrm{\gtrsim 10^{10}M_{\odot}}$). This relation has been studied for the local Universe and redshifts up to $z\sim 10$ (maiolino_amaze_2008; lamareille_physical_2009; mannucci_fundamental_2010; maier_mass-metallicity_2014; cresci_fundamental_2019; curti_mass-metallicity_2020; curti_chemical_2023; langeroodi_evolution_2023; nakajima_jwst_2023). The MZR has also been studied at spatially resolved scales (rosales-ortega_new_2012; sanchez_mass-metallicity_2013; sanchez_almeida_fundamental_2019; baker_metallicitys_2022). This relation has been observed to evolve with redshift: galaxies in the early universe are metal-poorer than local galaxies (tremonti_origin_2004; savaglio_gemini_2005; maiolino_amaze_2008; gao_what_2018; sanders_mosdef_2021). Generally, this is attributed to galaxies in the local universe having had more time to assemble and produce metals than at high redshifts.
@@
The second most widely studied relation is the metallicity’s anti-correlation with the star formation rate (SFR), which, together with the stellar mass, gives rise to the so-called Fundamental Metallicity Relation (FMR; mannucci_fundamental_2010; lara-lopez_fundamental_2010; cresci_fundamental_2019; curti_mass-metallicity_2020; pistis_comparative_2024; curti_jades_2024; sarkar_unveiling_2025). This relation has also been observed to exist at spatially resolved scales (baker_metallicitys_2022; koller_magpi_2024). Other significant relations of the gas-phase metallicity have been found with the stellar age (lian_massmetallicity_2015; sanchez-menguiano_local_2020), rotational velocity (garnett_luminosity-metallicity_2002; pilyugin_oxygen_2004; hughes_role_2013), galaxy size (ellison_clues_2008; sanchez_almeida_origin_2018), velocity dispersion (li_sdss-iv_2018; cappellari_full_2023), and gas mass fraction (bothwell_fundamental_2013; barrera-ballesteros_sdss-iv_2018; brown_role_2018).
@@
Nevertheless, stellar mass remains the strongest contributor to defining metallicity among those, and the origin of the MZR has long been debated (maiolino_re_2019). A common explanation is that massive galaxies inhabit a deeper gravitational potential that is more efficient at retaining metals against outflows compared to lower mass galaxies (tremonti_origin_2004; tumlinson_large_2011; chisholm_metal-enriched_2018). In this sense, both the dynamical mass ($\mathrm{M_{dyn}}$) and gravitational potential, defined as $\mathrm{\Phi_{dyn}=M_{dyn}/R_{e}}$, would form a stronger correlation with the metallicity than the stellar mass would. Dynamical measurements are much rarer than photometric or spectroscopic estimates; thus, a common approximation of the gravitational potential is $\mathrm{\Phi_{*}=M_{*}/R_{e}}$ (deugenio_gas-phase_2018; cappellari_full_2023).
@@
deugenio_gas-phase_2018 investigated the correlation between the metallicity and the stellar mass $\mathrm{M_{*}}$, gravitational potential $\mathrm{\Phi_{*}}$, and surface mass density $\mathrm{\Sigma_{*}=M_{*}/R_{e}^{2}}$ for a volume-limited sample from SDSS DR7. They found that the metallicity more tightly correlates with $\mathrm{\Phi_{*}}$ rather than $\mathrm{M_{*}}$ or $\mathrm{\Sigma_{*}}$.
@@
sanchez-menguiano_stellar_2024 used data from the MaNGA pyPipe3D value-added catalogue (VAC) of more than 3000 galaxies to simultaneously investigate the effect of 148 parameters on the gas-phase metallicity measured at $\rm 1R_{e}$ via random forest regression. Their results confirm that $\mathrm{\Phi_{*}}$, as a proxy for the baryonic gravitation potential, is the dominating property in determining the gas-phase metallicity. They also find that the photometric mass is more important in defining the metallicity than the stellar mass measured via spectroscopy.
boardman_tight_2024 likewise used MaNGA data and reported $\mathrm{\Phi_{*}}$ to correlate even more tightly with the gaseous $\rm N/O$ abundance ratio than with the gas-phase metallicity. boardman_competing_2025 subsequently found both abundance measures to possess secondary correlations with galaxies’ positions along the star-forming sequence, analogous to the FMR.
@@
A tight correlation with $\mathrm{\Phi_{*}}$ has also been found for the stellar metallicity (barone_sami_2018; vaughan_sami_2022; cappellari_full_2023; sanchez-menguiano_more_2024), consistent with the stellar fundamental metallicity relation reported by looser_stellar_2024, who found that $\mathrm{\Phi_{*}}$ provides a tighter correlation than stellar mass alone.
@@
On the contrary, an investigation done by baker_stellar_2023 found that the stellar mass $\mathrm{M_{*}}$, not $\mathrm{\Phi_{*}}$ or the dynamical measurements $\mathrm{M_{dyn}}$ and $\mathrm{\Phi_{dyn}}$, is the dominating factor in determining the gas-phase metallicity. Their study focused on Partial Correlation Coefficient (PCC) analysis and random forest regression (bluck_are_2020; baker_almaquest_2022) of roughly 1000 galaxies from the MaNGA survey.
@@
Throughout this paper, we assume a kroupa_variation_2001 initial mass function (IMF) and adopt a flat $\mathrm{\Lambda CDM}$ cosmology with $\mathrm{H_{0}=70.0kms^{-1}Mpc^{-1}}$, $\mathrm{\Omega_{m}=0.3}$, and $\mathrm{\Omega_{\lambda}=0.7}$.
@@ 2.1 MaNGA
We utilise publicly available integral field spectroscopic (IFS) data from MaNGA (Mapping Nearby Galaxies at Apache Point Observatory, bundy_overview_2015), part of the fourth-generation Sloan Digital Sky Survey (SDSS-IV, blanton_sloan_2017). We use the final and fully complete data release (DR) 17, released in 2022, for 10,000 galaxies (abdurrouf_seventeenth_2022; sanchez_sdss-iv_2022).
@@
We also use dynamical masses $\mathrm{M_{dyn}}$ and effective velocity dispersions, $\mathrm{\sigma_{e}}$, within elliptical half-light isophotes provided by zhu_manga_2023, which were computed via the axisymmetric Jeans Anisotropic Modelling (JAM, cappellari_measuring_2008; cappellari_efficient_2020).
@@ 2.2 SDSS
Our second dataset consists of data from the Sloan Digital Sky Survey (SDSS) DR7 (abazajian_seventh_2009), which includes spectroscopy for around 930,000 galaxies. We get stellar masses and SFRs from the MPA-JHU catalogue [^2] https://wwwmpa.mpa-garching.mpg.de/SDSS/DR7/ (kauffmann_host_2003; brinchmann_physical_2004; salim_uv_2007) and use the provided emission line fluxes for computing gas-phase metallicities. We match this data to a catalogue containing morphological parameters (simard_catalog_2011).
@@ 2.3 Sample selection
Regarding the SDSS data, we employ a signal-to-noise ratio (SNR) cut $\rm S/N>3$ for our lines of interest to ensure accurate emission line detection for calculating gas-phase metallicities. We also limit the sample by excluding highly inclined galaxies by applying a minor-to-major axis ratio cut of $\mathrm{b/a\geq 0.35}$. Furthermore, due to the small aperture of the SDSS observations, some galaxies which are at a very low redshift might only be probed within their most central region. To reduce aperture effects, we restrict our sample to galaxies for which the spectroscopic fibre radius covers more than $0.5\mathrm{R_{e}}$. The calibrations used for computing gas-phase metallicities are based on star-forming regions; thus, we select star-forming galaxies based on the $\mathrm{[NII]/H\alpha}$ vs. $\mathrm{[OIII]/H\beta}$ Baldwin-Philips-Terlevich (BPT) diagnostic diagram (baldwin_classification_1981), using the empirical line by kauffmann_host_2003, which separates star-forming galaxies from AGN/LINER galaxies. Our final SDSS sample consists of 121,880 star-forming galaxies.
@@ 2.4 Derivation of Physical Quantities
Stellar masses per pixel, integrated total stellar masses and masses within $\mathrm{1R_{e}}$ are obtained from pyPipe3D (sanchez_sdss-iv_2022). We obtain galactocentric distances for each spaxel in units of the effective radius from DAP (belfiore_data_2019; westfall_data_2019). Via the galactocentric distances, we compute total masses and masses within $\mathrm{1R_{e}}$ by taking the sum of all spaxels.
Emission lines are corrected for extinction using the $\mathrm{H\alpha/H\beta}$ Balmer decrement, assuming an intrinsic value of $2.86$ (osterbrock_astrophysics_2006), via the calzetti_dust_2000 dust attenuation law with $\mathrm{R_{v}=3.1}$.
Star formation rates for each spaxel in the MaNGA data are computed via the conversion law from kennicutt_star_2012 using the $\mathrm{H\alpha}$ emission line:
@@
We obtain a total SFR within $1\mathrm{R_{e}}$ by taking the sum of SFRs of the spaxels which fulfil our $\rm S/N$ and BPT diagram requirements. The global catalogue of pyPipe3D data available online [^3] https://data.sdss.org/datamodel/files/MANGA_PIPE3D/MANGADRP_VER/PIPE3D_VER/SDSS17Pipe3D.html also provides integrated SFRs using only the spaxels compatible with star formation. The global values are computed using the kennicutt_star_1998 calibration for a Salpeter IMF. To convert to the new calibration using a Kroupa IMF, we use the conversion given in kennicutt_star_2012:
@@
Table 1: Diagnostics used for computing gas-phase metallicities as presented in curti_mass-metallicity_2020.
@@
To obtain a singular value for the gas-phase metallicities, we compute the metallicities per spaxel using nine different strong-line calibrations (see Table 1) and take the average over all spaxels from our selection within each galaxy. The diagnostics $\mathrm{R23,R2,R3,O32,N2,S2,O3S2,O3N2,}$ and $\mathrm{R3S2}$ are used in combination with the empirical calibrations presented in curti_new_2017; curti_mass-metallicity_2020. Using several diagnostics simultaneously decreases the risk of having degeneracies associated with diagnostics that are double values. It also allows for integrating information from several emission lines that may differ across metallicity regimes, and it allows us to constrain the resulting abundances more precisely (see curti_mass-metallicity_2020 for more information). We take the average metallicity value across all selected spaxels within $1\mathrm{R_{e}}$ to obtain a single value for the gas-phase metallicity per galaxy. We also repeated our analysis using the median and found no significant difference in the results.